OpenStudy (anonymous):

Can someone remind me what scientific notation is and can you provide an example ?

OpenStudy (anonymous):

1000= 1*10^3

OpenStudy (anonymous):

0.001= 1*10^-3

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

Scientific notation is a way of expressing very large or very small numbers without all the pesky zeros. For instance, 1 billion is 1,000,000,000 = 10^9. If you wanted to express 4 billion, it would be 4 x 10^9. It's used everyday without people realizing it. How many times have you heard someone talk about an athlete making $2.6 million (that's $2.6 x 10^6).

OpenStudy (anonymous):

Scientific notation, also known as standard form or as exponential notation, is a way of writing numbers that accommodates values too large or small to be conveniently written in standard decimal notation. Scientific notation has a number of useful properties and is commonly used in calculators, and by scientists, mathematicians, doctors, and engineers.